# Graph Neural Network Enhanced Retrieval for Question Answering of LLMs

### v1

@@ 3.4 RGNN Enhanced Retrieval for Multi-hop Reasoning Questions
The information asymmetry phenomenon also frequently arises in complex multi-hop reasoning questions, as analyzed in the Introduction section. Although recent works have incorporated multi-hop reasoning or question rewriting into retrieval [35, 30, 27], LLMs may generate hallucinated reasons or incorrect subquestions in the absence of prior knowledge in this domain [7]. These hallucinated reasons or incorrect subquestions fail to retrieve supporting passages when only considering the retrieval process in isolation for each step. Additionally, they also encounter another challenge: whether to continue the next-step answering process or output the final answer. To overcome these challenges, we first propose a self-critique technique by prompting LLMs to determine the termination of the answering process. To enhance retrieval of supporting passages over steps, we further utilize an RGNN to enable the integration between graphs of passages in different steps.
@@
Self-critique of LLMs. To answer multi-hop reasoning questions, self-critique involves prompting the LLM to generate subquestions and answer them using retrieved passages in a step-by-step manner. As depicted in Fig. 3, at each step, after the LLM generates a subanswer to the subquestion, the question and all the generated subanswers are input into the LLM and critiqued to determine whether the generated subanswers are evident to generate the final answer to the question. If the output is ‘Yes’, the LLM will be prompted to generate the final answer. Otherwise, it will be requested to generate the next-step subquestion. Compared with the prior work, SelfAsk [23], we focus on whether existing evidences is satisfactory for answering the question instead of whether the next-step subquestion is needed. This self-critique technique decomposes the question over steps and enables a more accurate answer to the question.
@@
Recurrent graph neural network. To enhance retrievals over steps, as shown in Fig. 3, we utilize an RGNN to establish relationships between the semantic distances of subquestions. Specifically, with the subquestion $q^{t}$ at step $t$, we first compute $h_{i}^{L,t}$ for each node $v_{i}\in\mathcal{V}$ using the aforementioned $L$-layer GNN with Eq. (3) and (4). To consider the effect of semantic distances from previous subquestions, we integrate $h_{i}^{L,t}$ with the previous-step integrated semantic distance to compute the integrated semantic distance $\hat{h}_{i}^{L,t}$ at step $t$ using the parameter $\beta$. This can be formulated as follows:
$\hat{h}_{i}^{L,t}=\beta h_{i}^{L,t}+(1-\beta)\hat{h}_{i}^{L,t-1},\quad(t>0),$ (6)
@@
where $\hat{h}_{i}^{L,t-1}$ is the integrated semantic distance from the previous step $t-1$ and $\hat{h}_{i}^{L,0}=h_{i}^{L,0}$. The integrated semantic distance $\hat{h}_{i}^{L,t}$ will be used to integrate with ${h}_{i}^{L,t+1}$ of node $v_{i}$ at the next step $t+1$ recurrently until the last step of the answering process $T$. GNN facilitates the integration of semantic distances between related passages at each step, and the recurrent parameter $\beta$ enables the integration of semantic distances of different subquestions across steps, which effectively mitigates the impacts of incorrect subquestions and enhances the retrieval of supporting passages for them.
@@
$\ell$ $=\frac{1}{T}\sum_{t\in[T]}\ell^{t}=\frac{1}{T}\sum_{t\in[T]}\max(0,r+\bar{d}_{Y^{t+}}^{L,t}-\bar{d}_{O/Y^{t+}}^{L,t}).$ (7)
By employing this hinge objective, the RGNN takes into account the retrieval of subsequent answering steps and proactively reduces the semantic distances associated with them. These small semantic distances are then transferred to the next subquestion and enhance the retrieval of the corresponding supporting passages. We update the RGNN (parameters $\alpha^{l}$ and $\beta$) using the gradient descend, and more training details are elaborated in Appendix A.2.
@@ 4 Experiments
In this section, we aim to conduct experiments to demonstrate the effectiveness of GNN-Ret and RGNN-Ret as follows: 1) Our proposed GNN-Ret and RGNN-Ret achieve superior accuracy on various QA datasets (Table 1). 2) We conduct ablation studies to select hyperparameters $K$ and $O$ for training GNN and RGNN (Fig. 4) and verify the effectiveness of each component in GNN-Ret and RGNN-Ret (Table 2). 3) GNN-Ret improves the retrieval coverage of supporting passages (Fig. 5).
@@ 4.1 Experimental Setups
Evaluation datasets. We measure all the methods on four different QA datasets, including multi-hop Wikipedia reasoning datasets: (1) MuSiQue [29], (2) IIRC [4], (3) 2WikiMQA [6], and a single-hop multi-choice QA dataset: (4) Quality [21]. For each multi-hop dataset, we randomly sample 500 questions from the development set, which provides the ground-truth supporting passages. We use 20 of them to train the GNN and select the rest as test data. We use all the Wikipedia documents of these 500 questions as the retrieval corpus, which is significantly larger than that provided by datasets [^1] Prior works often use 10–20 paragraphs or documents as the retrieval corpus for each question [27, 33].. For each document or article, we split it into multiple passages with a maximum token length of 200 [25]. As evaluation metrics, we calculate the F1 score, exact match (EM) and accuracy (Acc) for multi-hop reasoning datasets. We use the ChatGPT (gpt-3.5-turbo-2023-06-01-preview) to evaluate if the prediction matches with the gold answer. For Quality, we randomly sample 30 articles with overall corresponding 561 questions and use these articles as the retrieval corpus. Since Quality is not a multi-hop dataset, we only validate the accuracy performance of GNN-Ret on it.
Baselines. We compare GNN-Ret and RGNN-Ret with the following baselines: (1) Direct answers questions without retrieved passages. (2) Dense [14] answers questions with retrieved passages using dense retrieval. (3) SelfAsk [23] prompts LLMs to generate the follow-up question and answers it with retrieved passages until generating the final answer. (4) ITER-RETGEN [27] iteratively answers questions with retrieved passages and uses the generated answer for the next-step retrieval until the generation of final answer. (5) IRCoT [30] iteratively prompts LLMs to generate chains of thoughts with retrieved passages and retrieves with the generated reasons until reaching the maximum token number. All the retrieved passages are then used to generate the final answer. (6) KGP [33] first searches seed passages using the question and then prompts LLMs to generate the needed evidence to retrieve the other relevant passages among neighbors of seed nodes using semantic distances.
Implementation details. We use the ChatGPT (gpt-3.5-turbo-2023-06-01-preview) as the LLM backbone for experiments and adopt a temperature of 0 to remove the effect of random sampling. We select mpnet-base-cos-v1222https://huggingface.co/sentence-transformers/multi-qa-mpnet-base-cos-v1 as the embedding model to compute the semantic distance $=1-\text{cosine similarity}$ following [25] for all the methods. We set up a maximum token number of 3500 for retrieval to leave the space to instruction and demonstrations, and all the methods retrieve the semantically similar passages until reaching the maximum token number. We adopt a one-layer GNN for our proposed methods GNN-Ret and RGNN-Ret. We set up $\alpha^{1}=0.5$ for GNN-Ret and $\alpha^{1}=0.5$ and $\beta=0.9$ for RGNN-Ret when there is no training. When training the RGNN, we set up $r=0.01$, $K=5$, and $O=25$ for GNN-Ret (w. train) and $r=0.01$, $K=5$, and $O=10$ for RGNN-Ret (w. train), respectively, according to the ablation studies in Fig. 4. We do not train the parameters for the Quality dataset without the ground-truth labels of supporting passages. Since training the RGNN requires ground-truth subquestions, we manually generate subquestions for 5 questions sampled from the preset training data. More implementation details are present in Appendix B.
@@
SelfAsk [23] 28.0 18.3 22.9 37.1 29.6 36.9 51.2 41.5 44.8 - - -
ITER-RETGEN [27] 30.0 16.7 27.1 32.6 20.0 39.6 45.6 33.3 43.8 - - -
IRCoT [30] 29.9 12.5 27.5 32.5 18.8 41.0 46.0 30.8 44.6 - - -
KGP [33] 30.0 15.4 24.0 33.9 18.5 42.1 45.5 31.9 45.2 - - -
@@ 4.2 Main Results
Table 1 summarizes the results for our proposed methods and baselines on four QA datasets using ChatGPT. The results show that our proposed GNN-Ret significantly outperforms baselines Dense and Direct in terms of F1, EM, and accuracy on all these four QA datasets with a single retrieval. For instance, compared with Dense, GNN-Ret improves EM by 4.4 and accuracy by $4\%$ on the IIRC dataset. Additionally, GNN-Ret maintains its superiority on the Quality dataset, which requires a comprehensive understanding of the context of an entire story. Notably, GNN-Ret effectively enhances accuracy by $4.8\%$ with $K=10$ on the Quality dataset. This improvement can be attributed to the ability of GNN-Ret to enhance the retrieval prioritization of supporting passages through the incorporation of structural information, thereby improving the comprehension of LLMs in articles. Surprisingly, our proposed one-hop GNN-Ret even slightly outperforms baselines that perform multi-hop retrieval processes and queries of LLMs in accuracy, which highlights the significant potential of leveraging passage relatedness to enhance the retrieval process for answering multi-hop reasoning questions. Furthermore, the results in Table 1 demonstrate that our proposed RGNN-Ret achieves state-of-the-art performance in terms of F1, EM, and accuracy on the three multi-hop reasoning datasets. RGNN-Ret outperforms the best baseline, KGP, by more than $10\%$ in accuracy on 2WikiMQA dataset. This performance improvement can be attributed to our proposed self-critique technique and RGNN. The self-critique technique enables more accurate judgement on the ending of the answering process and the generation of final answers, and the incorporation of RGNN further enhances the retrieval process for each step of answering, which is essential for generating subquestions and subanswers.
@@ 4.3 Additional Analysis
Selection of $K$ and $O$. In order to effectively train the GNN and RGNN models, we conduct ablation studies to select hyperparameters $K$ and $O$, which determine the number of relevant nodes to be sampled and the size of the competitive node set, respectively. To assess the impact of different values of $K$ and $O$, we evaluate the accuracy of GNN-Ret and RGNN-Ret on the MuSiQue and IIRC datasets with varying $K$ and $O$. The average accuracy across these two datasets is represented by the green points in Fig. 4. The results demonstrate that both GNN-Ret and RGNN-Ret consistently achieve stable accuracy performance across different settings of $K$ and $O$. Upon analyzing the results, we observe that GNN-Ret achieves the highest accuracy when $K=5$ and $O=25$, while RGNN-Ret performs best with $K=5$ and $O=10$. GNN-Ret adopts a larger value of $O$ since it considers the whole label set during training, while RGNN-Ret only considers the labels for the current step and subsequent steps. Therefore, we select these settings of $K$ and $O$ to train the GNN and RGNN models for experiments reported in Table 1.
(a) GNN-Ret
(b) RGNN-Ret
@@
Effectiveness of components in RGNN-Ret. We conduct ablation studies to evaluate the effectiveness of each component in RGNN-Ret, and present results in Table 2. We first remove the RGNN and solely performs dense retrieval for our proposed self-critique technique. The results in Table 2 show that it results in an accuracy reduction of $2.1\%$ and $2.7\%$ on MuSiQue and 2WikiMQA datasets, respectively. Next, we incorporate the GNN for the retrieval process of each subquestion, i.e., Self-critique + GNN. This modification improves accuracy on both MuSiQue and 2WikiMQA datasets compared to the settings with dense retrieval, indicating the effectiveness of the GNN in enhancing the retrieval coverage for each subquestion individually. By combining self-critique and RGNN, our proposed method RGNN-Ret achieves accuracy up to $31.3\%$ and $55.6\%$ on MuSiQue and 2WikiMQA datasets, respectively, which demonstrates the effectiveness of RGNN-Ret in enhancing the retrieval by interconnecting with the graph of passages from previous steps. This interconnection enhances the retrieval of supporting passages across multiple steps, leading to more accurate answers in the multi-hop reasoning process.
Method MuSiQue 2WikiMQA
Dense 23.8 42.3
GNN-Ret (mean) 26.3 44.6
GNN-Ret 27.3 46.9
Self-critique + dense 29.2 52.9
Self-critique + GNN 30.8 55.2
RGNN-Ret 31.3 55.6
@@
RGNN-Ret ($Y^{t}$) 30.2 54.6
RGNN-Ret 31.3 55.6
Table 2: Ablation studies of components of GNN and RGNN in accuracy with ChatGPT on MuSiQue and 2WikiMQA datasets.
@@
Statistics of retrieval accuracy. We explore the EM between the retrieved passages and the ground-truth passages and collect the exact-match number of test samples with varying numbers of supporting passages required in questions on MuSiQue and 2WikiMQA datasets. We do not evaluate the statistics of retrieval accuracy for RGNN-Ret without the specific ground-truth supporting passages for each subquestion. The results in Fig. 5 show that our proposed GNN-Ret achieves higher retrieval accuracy compared with Dense on MuSiQue and 2WikiMQA datasets. For instance, GNN-Ret outperforms Dense by 29 and 43 exact-match test samples for the questions that require 2 supporting passages. This demonstrates that the GNN indeed improves retrieval coverage of supporting passages and the retrieval prioritization of supporting passages for subsequent reasoning. We display some case studies to demonstrate advantages of GNN-Ret and RGNN-Ret in Appendix C.
(a) MuSiQue
(b) 2WikiMQA

### v2

@@ 3.3 RGNN Enhanced Retrieval for Multi-hop Reasoning Questions
The information asymmetry phenomenon also frequently arises in complex multi-hop reasoning questions, as analyzed in the Introduction section. Although recent works have incorporated multi-hop reasoning or question rewriting into retrieval Yao et al. (2022); Trivedi et al. (2023); Shao et al. (2023), LLMs may generate hallucinated reasons or incorrect subquestions due to the absence of prior knowledge to the question Huang et al. (2023). These hallucinated reasons or incorrect subquestions fail to retrieve supporting passages when only considering the retrieval process in isolation for each step. Additionally, they also encounter another challenge: whether to continue the next-step answering process or output the final answer. To overcome these challenges, we first propose a self-critique technique by prompting LLMs to determine the termination of the answering process. To enhance retrieval of supporting passages over steps, we further utilize an RGNN to enable the integration between graphs of passages in different steps.
@@
Self-critique of LLMs. To answer multi-hop reasoning questions, self-critique involves prompting the LLM to generate subquestions and answer them using retrieved passages in a step-by-step manner. As depicted in Fig. 3, at each step, after the LLM generates a subanswer to the subquestion, the question and all the generated subanswers are input into the LLM and critiqued to determine whether the generated subanswers are evident to generate the final answer to the question. If the output is ‘Yes’, the LLM will be prompted to generate the final answer. Otherwise, it will be requested to generate the next-step subquestion. Compared with SelfAsk Press et al. (2022), we omit the subquestions and focus on whether existing evidences are satisfactory for answering the question when determining the termination. This self-critique technique decomposes the question over steps and enables a more accurate answer to the question.
@@
Recurrent graph neural network. To enhance retrievals over steps, as shown in Fig. 3, we utilize an RGNN to establish relationships between the semantic distances of subquestions. Specifically, with the subquestion $q^{t}$ at step $t$, we first compute $h_{i}^{L,t}$ for each node $v_{i}\in\mathcal{V}$ using the aforementioned $L$-layer GNN. To consider the effect of semantic distances from previous subquestions, we integrate $h_{i}^{L,t}$ with the previous-step integrated semantic distance to compute the integrated semantic distance $\hat{h}_{i}^{L,t}$ at step $t$ using the parameter $\beta$. This can be formulated as follows:
$\hat{h}_{i}^{L,t}=\beta h_{i}^{L,t}+(1-\beta)\hat{h}_{i}^{L,t-1},\quad(t>1),$ (4)
@@
where $\hat{h}_{i}^{L,t-1}$ is the integrated semantic distance from the previous step $t-1$ and $\hat{h}_{i}^{L,1}=h_{i}^{L,1}$. The integrated semantic distance $\hat{h}_{i}^{L,t}$ will be used to integrate with ${h}_{i}^{L,t+1}$ of node $v_{i}$ at the next step $t+1$ recurrently until the last step of the answering process $T$. GNN facilitates the integration of semantic distances between related passages at each step, and the recurrent parameter $\beta$ enables the integration of semantic distances of different subquestions across steps, which effectively mitigates the impacts of incorrect subquestions and enhances the retrieval of supporting passages for them.
@@
$\ell=\frac{1}{T}\sum_{t\in[T]}\max(0,r+\bar{d}_{Y^{t+}}^{L,t}-\bar{d}_{O/Y^{t+}}^{L,t}).$ (5)
By employing this hinge objective, the RGNN takes into account the retrieval of subsequent answering steps and proactively reduces the semantic distances associated with them. These small semantic distances are then transferred to the next subquestion and enhance the retrieval of the corresponding supporting passages. We update the RGNN (parameters $\alpha^{l}$ and $\beta$) using the gradient descend, which are elaborated in Appendix A.2.
@@ 4 Experiments
In this section, we conduct experiments to demonstrate the effectiveness of GNN-Ret and RGNN-Ret, including 1) superiority of GNN-Ret and RGNN-Ret on various QA datasets (Table 1), 2) ablation studies to verify the effectiveness of components in GoPs, GNN-Ret, and RGNN-Ret (Table 2 and 3), and 3) superiority of GNN-Ret on open-sourced LMs (Table 4). Additional analysis on hyperparameter selection, statistics of retrieval accuracy, token number for retrieval, multi-layer GNN-Ret, extension of RGNN-Ret to other multi-hop answering methods, and qualitative results are discussed in Appendix C, D, E, F, G, H, respectively.
@@ 4.1 Experimental Setups
Evaluation datasets. We measure all the methods on four different QA datasets, including multi-hop Wikipedia reasoning datasets: (1) MuSiQue Trivedi et al. (2022), (2) IIRC Ferguson et al. (2020), (3) 2WikiMQA Ho et al. (2020), and a single-hop multi-choice QA dataset: (4) Quality Pang et al. (2021). As evaluation metrics, we calculate the F1 score, exact match (EM) and accuracy (Acc) for multi-hop reasoning datasets. We use the ChatGPT (gpt-3.5-turbo-2023-06-01-preview) to evaluate if the prediction matches with the gold answer, following Shao et al. (2023); Wang et al. (2024). Since Quality is not a multi-hop dataset, we only validate accuracy performance of GNN-Ret on it.
Baselines. We compare GNN-Ret and RGNN-Ret with the following baselines: (1) Direct answers questions without retrieved passages. (2)We use retrievers bm25 Robertson et al. (2009), DPR Karpukhin et al. (2020a), and SentenceBert [^1] https://huggingface.co/sentence-transformers/multi-qa-mpnet-base-cos-v1 (SBERT) Reimers and Gurevych (2019) to retrieve passages without considering relatedness between passages for QA. (3) SelfAsk Press et al. (2022) prompts LLMs to generate the follow-up question and answers it with retrieved passages until generating the final answer. (4) ITER-RETGEN Shao et al. (2023) iteratively answers questions with retrieved passages and uses the generated answer for the next-step retrieval until the generation of final answer. (5) IRCoT Trivedi et al. (2023) iteratively prompts LLMs to generate chains of thoughts with retrieved passages and retrieves with the generated reasons until reaching the maximum token number. All the retrieved passages are then used to generate the final answer. (6) KGP Wang et al. (2024) first searches seed passages using the question and then prompts LLMs to generate the needed evidence to retrieve the other relevant passages among neighbors of seed nodes using semantic distances.
@@
SelfAsk 28.0 18.3 22.9 37.1 29.6 36.9 51.2 41.5 44.8 - - -
ITER-RETGEN 30.0 16.7 27.1 32.6 20.0 39.6 45.6 33.3 43.8 - - -
IRCoT 29.9 12.5 27.5 32.5 18.8 41.0 46.0 30.8 44.6 - - -
KGP 30.0 15.4 24.0 33.9 18.5 42.1 45.5 31.9 45.2 - - -
@@ 4.2 Main Results
Table 1 summarizes the results for our proposed methods and baselines on four QA datasets using ChatGPT. Results show that GNN-Ret significantly outperforms baselines in terms of F1, EM, and accuracy on all these four QA datasets with a single retrieval. For instance, compared with SBERT, GNN-Ret improves EM by 4.4 and accuracy by $4\%$ on the IIRC dataset. Additionally, GNN-Ret maintains its superiority on the Quality dataset, which requires a comprehensive understanding of the context of an entire story. This can be attributed to the ability of GNN-Ret to enhance the retrieval prioritization of supporting passages through the incorporation of structural information. Surprisingly, our proposed one-hop GNN-Ret even slightly outperforms baselines with multi-hop retrieval processes and queries of LLMs in accuracy, highlighting significant potentials of leveraging passage relatedness to enhance the retrieval process for answering multi-hop reasoning questions.
Furthermore, results in Table 1 demonstrate that our proposed RGNN-Ret achieves state-of-the-art performance in terms of F1, EM, and accuracy on the three multi-hop reasoning datasets. RGNN-Ret outperforms the best baseline, KGP, by more than $10\%$ in accuracy on 2WikiMQA. This performance improvement can be attributed to our proposed self-critique technique and RGNN. The self-critique technique enables more accurate judgement on the ending of the answering process and the generation of final answers. The incorporation of RGNN further enhances retrieval for each step of answering.
@@ 4.3 Additional Analysis
Effectiveness of components in RGNN-Ret. We conduct ablation studies to evaluate the effectiveness of each component in RGNN-Ret (i.e., recurrent, GNN, and RGNN). Results in Table 2 show that both the recurrent part (parameter $\beta$) and GNN improve accuracy compared with setting with SBERT on MuSiQue and 2WikiMQA datasets. By combining the recurrent part and GNN, our proposed method RGNN-Ret further improves accuracy up to $31.3\%$ and $55.6\%$ on MuSiQue and 2WikiMQA datasets, respectively, which demonstrates the effectiveness of RGNN-Ret in enhancing the retrieval by interconnecting with the GoPs from previous steps. This interconnection enhances the retrieval of supporting passages across multiple steps, leading to more accurate answers in the multi-hop reasoning process.
Methods MuSiQue 2WikiMQA
SBERT 23.8 42.3
GNN-Ret (mean) 26.3 (+2.5) 44.6 (+1.7)
GNN-Ret 27.3 (+3.5) 46.9 (+4.6)
Self-critique + SBERT 27.9 52.9
Self-critique + recurrent 28.1 (+0.2) 54.0 (+1.1)
Self-critique + GNN 30.8 (+2.9) 55.2 (+2.3)
RGNN-Ret 31.3 (+3.4) 55.6 (+2.7)
@@
RGNN-Ret ($Y^{t}$) 30.2 (+3.1) 54.6 (+1.3)
RGNN-Ret ($Y^{t+}$) 31.3 (+4.2) 55.6 (+2.3)
Table 2: Ablation studies of components of GNN-Ret and RGNN-Ret in accuracy with ChatGPT on MuSiQue and 2WikiMQA.
@@
Effectiveness of GoPs. We explore the effectiveness of components (structural information (SI) and shared keywords (SK)) for graph construction. Results in Table 3 show that both structural information and shared keywords are able to improve accuracy for the baseline without GoPs. Notably, in the context-specific Quality story dataset, which demands a comprehensive grasp of the entire narrative for accurate retrieval, the graph that incorporates structural information markedly outperforms the setting that relies solely on shared keywords. While for 2WikiMQA that requires knowledge of multiple documents, the graph with shared keywords achieves a better accuracy compared with that with structural information. By combining them into graph construction, we achieve the best accuracy in these two datasets.
Methods 2WikiMQA Quality
SBERT (w.o. graph) 42.3 53.4
graph w. SI 42.7 (+0.4) 55.5 (+2.1)
graph w. SK 45.4 (+3.1) 54.8 (+1.4)
graph w. SI and SK 46.9 (+4.6) 58.2 (+4.8)
Table 3: Ablation studies of components on graph construction in accuracy with ChatGPT on 2WikiMQA and Quality. SI and SK represent structural information and shared keywords, respectively.
Performance of GNN-Ret with open-sourced LMs. Our proposed GNN-Ret utilizes the relatedness of passages to enhance retrieval with the meticulously designed GNNs, which is generalized and robust to all the LM architectures. To substantiate this, we supplement the experiments for GNN-Ret and SBERT with two well-behaved open-sourced LMs (Qwen/Qwen2-7B-Instruct Yang et al. (2024) and google/gemma-2-9b-it Team (2024)) from Open LLM Leaderboard Fourrier et al. (2024). Results in Table 4 show that GNN-Ret consistently outperforms SBERT across varying open-sourced LMs, demonstrating the generalization of GNN-Ret in enhancing retrieval coverage with varying LMs.
MuSiQue 2WikiMQA
Qwen2-7B-Instruct F1 EM Acc F1 EM Acc
SBERT 27.8 16.9 23.3 43.3 34.0 40.8
GNN-Ret 32.3 20.8 26.7 50.5 37.3 45.8
gemma-2-9b-it F1 EM Acc F1 EM Acc
SBERT 32.8 18.9 26.9 46.4 28.8 43.8
GNN-Ret 34.6 20.8 29.0 51.6 30.8 45.8
Table 4: F1/EM/Acc of SBERT and GNN-Ret with open-sourced LMs on MuSiQue and 2WikiMQA.